# A Bio-Inspired Research Paradigm of Collision Perception Neurons Enabling Neuro-Robotic Integration: The LGMD Case

### v1

@@
1cm1cm Compared to human vision, insect visual systems excel at rapid and precise collision detection, despite relying on only tens of thousands of neurons organized through a few neuropils. This efficiency makes them an attractive model system for developing artificial collision-detecting systems. Specifically, researchers have identified collision-selective neurons in the locust’s optic lobe, called lobula giant movement detectors (LGMDs), which respond specifically to approaching objects.
Research upon LGMD neurons began in the early 1970s. Initially, due to their large size, these neurons were identified as motion detectors, but their role as looming detectors was recognized over time. Since then, progress in neuroscience, computational modeling of LGMD’s visual neural circuits, and LGMD-based robotics has advanced in tandem, each field supporting and driving the others. Today, with a deeper understanding of LGMD neurons, LGMD-based models have significantly improved collision-free navigation in mobile robots including ground and aerial robots.
This review highlights recent developments in LGMD research from the perspectives of neuroscience, computational modeling, and robotics. It emphasizes a biologically plausible research paradigm, where insights from neuroscience inform real-world applications, which would in turn validate and advance neuroscience. With strong support from extensive research and growing application demand, this paradigm has reached a mature stage and demonstrates versatility across different areas of neuroscience research, thereby enhancing our understanding of the interconnections between neuroscience, computational modeling, and robotics. Furthermore, other motion-sensitive neurons have also shown promising potential for adopting this research paradigm.
@@ 1 Introduction
Visually guided collision avoidance is a common behavior in many sighted animal species. Unlike humans, insects can accurately perceive impending collisions with only a few thousand neurons, making them ideal models for designing efficient and simplified collision detection systems.
@@
Anatomical research on the LGMDs and circuitry was inaugurated in the 1970s [3, 4, 5, 6], while two distinct types of computational models of LGMD1 emerged in the 1990s [7, 8]. The first type of computational models incorporates physical attributes, i.e., image size and speed, to represent the LGMD’s response to looming objects, focusing on dendritic computations within the neuron. In contrast, the second type of models utilizes critical image cues and lateral inhibition within the signal processing pathway to LGMD1 to simulate its neural responses, enabling the distinction between approaching, receding, and translating stimuli. Although LGMD2 shows the similar looming selectivity222neuron responding most strongly to approaching rather than other kinds of visual movements of LGMD1, computational modeling of LGMD2 has emerged late in recent years due to very limited biological findings, compared to LGMD1 [9].
In real-world scenarios, the size and speed of an approaching object are often unknown, unable to measure, which poses significant challenges for implementing the first type of computational model in complex robot scenarios interacting with dynamic environments. Moreover, this model only predicts the LGMD1 response to approaching stimuli, without accounting for other types of motion like receding and translating. In contrast, the second type of model has been successfully extended and demonstrated in real-world environments using a mobile robot system [10], marking a significant milestone in the development of LGMD-based robotics.
@@ 2 How LGMD works as looming detectors
2.1 Classic biological and computational theory on LGMD
@@
Figure 2: The morphology of LGMD1 and LGMD2 with response to different visual stimuli. (A) The afferent network of LGMD1 (adapted from [26]). LGMD1 is located in the fourth layer of the optic lobe in locusts, with the retina, lamina, and medulla serving as its afferent layers. (B) The morphology of LGMD1 (adapted from [27]). LGMD1 has three distinct dendritic fields where field A receives approximately $\sim 15,000$ excitatory retinotopic inputs from the entire visual hemifield, while fields B and C receive approximately $\sim 500$ non-retinotopic feedforward inhibitory inputs related to ON and OFF contrasts, respectively. Recent research has shown that dendritic field C also contributes to processing excitatory ON-contrast signals. (C) Schematic diagram of LGMD1 and LGMD2 morphology. The LGMD2 neuron (illustrated in blue) has only a single large dendrite field, and its downstream signaling pathway and postsynaptic targets remain unclear. In contrast, the downstream signaling pathway of LGMD1 has been identified as the DCMD. Dendrite field A (depicted in green) receives excitatory inputs and encodes angular velocity in the OFF contrast. The “-” symbol indicates that lateral inhibition may also occur on the surface of the LGMD1. Dendrite fields B and C (shown in red) receive inhibitory signals from the ON and OFF pathways, respectively. Additionally, recent studies suggest that dendrite field C also processes non-retinotopic excitatory signals from the ON pathway. (D) The response of LGMD1 when a locust views a square projected on a screen approaching and then receding at $5m/s$ (adapted from [28]). LGMD1 shows a strong and continuous spike train as the square approaches, whereas the neuron displays only a phasic spike at the onset of receding. (E) The response of LGMD1 to approaching and translating stimuli (adapted from [19]) - regardless of the size and speed of the translating bar, the instantaneous firing rate of LGMD1 is significantly lower when the locust views translating stimuli compared to looming stimuli. (F) The neuronal responses of LGMD1 and LGMD2 when the locust views light and dark rectangles approaching and receding, as well as in response to changes in luminance (adapted from [9]) - LGMD2 is selectively excited by darker objects approaching. The long arrow indicates LGMD2 excitation after the onset of light object receding, while the short arrow indicates hyperpolarization of LGMD2 when both light and dark objects ceased during approach or at the start of receding of a dark object.
@@
Regarding computational modeling of LGMDs, the models of LGMD1 emerged since the 1990s [8], while the first model of LGMD2 was proposed over two decades later [31, 32]. In general, bio-inspired looming perception visual systems can be categorized into two primary classes. The first class focuses on mapping nonlinear relationships between the physical attributes of stimuli (image size and image velocity) and corresponding neuronal activity. The second class utilizes hierarchical neural networks modeled after well-established anatomical pathways responsible for looming perception. Assuming that dendritic field A of the LGMD1 neuron encodes angular velocity, while dendritic fields B and C encode angular size, Hatsopoulos et al. proposed the first computational model of LGMD1, known as the $\eta$-function [7]. As illustrated in Fig. 3A, for a looming object with half-size $l$ and constant approach speed $v$, the angular size $\theta$ projected on the retina can be described by $\tan(\theta)=\frac{2l}{v\cdot t}$, which increases rapidly as collision becomes imminent. The angular velocity $\theta^{\prime}(t)$ is the derivative of angular size with respect to time. The neuronal response of LGMD1 to looming stimuli, from a single neuron computational perspective, is modeled by the following $\eta$-function [^3] Responses of looming sensitive neurons in both vertebrates and invertebrates could also encode other optical variables like $\tau$ relative rate of expansion, $\rho$ absolute rate of expansion [33].:
@@
Researchers have demonstrated that the multiplicative combination of neuronal signals encoding an object’s angular size and angular speed can be achieved through a logarithmic-exponential transformation in the LGMD1 neuron [38], providing further evidence that the mathematically straightforward $\eta$-function can serves as the intrinsic computational mechanism for the neuronal response of LGMD1. However, the $\eta$-function models behave like the LGMD1 only for objects approaching at a constant speed [39], lacking the capacity to simulate detailed responses to accelerating objects and other types of visual stimuli, such as recession and translation.
On the other hand, Rind and Simmons proposed a four-layered neural network to model the LGMD1’s neuronal response to looming, receding, and translating stimuli [8] (see Fig. 3C). In this network, critical image cues (denoted as P in Fig. 3C) are generated by edges that change in extent or velocity during object approach, resulting in strong excitation of computational units (denoted as E in Fig. 3C). Lateral inhibition within these units is propagated with a time delay (denoted as I in Fig. 3C). An additional unit, termed the F unit (denoted as F in Fig. 3C), is integrated into the network to simulate the feed-forward inhibition (FFI) observed in the visual pathway to LGMD1. The captured critical image cues, the critical race between excitation and inhibition and the FFI together fine tune the looming selecticvity of the proposed neural network, showing strong and increasing response to looming stimuli, while transient response to object recession and minor response for single edge crossing (elongation) (see Fig. 3D, E). Reducing the strength of lateral inhibition significantly raises the network’s response to object receding and translating. Moreover, removing the FFI prolongs the network’s response to both object approach and recession.
@@ 2.2 Controversial arguments on LGMD neural processing
Although both categories of computational models effectively capture the neuronal characteristics of the LGMD1 when faced with looming objects, a key point of contention between these models lies in the timing of the peak response. In the first class of models, the $\eta$-function produces a peak response after a fixed delay, following the point at which the looming object reaches a specific image size threshold. However, the peak will be occurring after the object projects its maximum size onto the retina when $l/|v|\leq 5\text{ ms}$ [41]. This ratio can be achieved by, for example, an 8 cm-wide object moving at $8$ m/s ($28.8$ km/h) or a 4 cm-wide object traveling at $4$ m/s ($14.4$ km/h). In contrast, experimental results for the second category of models indicate that the peak response consistently occurs before the looming object reaches its maximum size on the retina.
@@
Secondly, while ON and OFF contrasts are separated in lamina monopolar cells (LMCs) located in the lamina, the precise mechanisms of how these signals combine or interact, particularly in the optic lobe of locusts, are unverified. Earlier research on LGMD1 suggested that ON and OFF contrast signals jointly flow into the medulla layer, where they transmit to dendritic field A for excitation, and to dendritic trees B and C for ON/OFF-contrast FFI, respectively [37]. However, recent neuroscience findings indicate that dendritic field C may partially contribute to ON contrast excitation, receiving inputs from a few neurons in the dorsal uncrossed bundle (DUB) neurons with large receptive fields (approximately $50\degree$). Meanwhile, other neurons with smaller receptive fields (approximately $10\degree$) are more likely responsible for the OFF-contrast FFI [16, 33]. Moreover, it has been demonstrated that global inhibition occurs in the lamina layer, functioning as a normalization mechanism for increasing excitation as collision approaches [49]. Computational modelers have attempted to replicate this function by introducing a non-spiking neuron in the lamina layer [22]. However, anatomical evidence may conflict with this modeling assumption, as each ommatidium has its own set of photoreceptors and LMCs, and LMCs are believed to interact only within their own set and not across different ommatidia. Nevertheless, the possibility of global inhibition in the lamina layer cannot be ruled out, as reconstructions suggest that many photoreceptors synapses in locusts involve four postsynaptic cells [50].
@@
Thirdly, regarding the trans-medullary afferent neurons (TmAs) as inputs to LGMDs (targeting dendritic field A of both LGMD1 and LGMD2), early neural modeling efforts assumed that TmAs were responsible for lateral inhibition in the LGMD pathway [21, 22]. When traced for reconstruction, TmAs indeed exhibited lateral connections and received input from the outer medulla layer. However, evidence suggests that this lateral interaction contributes to excitation rather than inhibition. It is, instead, in the lobula layer - at the surface of the LGMD - where lateral inhibition occurs [17, 51, 52, 49]. The lateral excitation involving TmAs is thought to enhance LGMD’s responses to coherent stimuli [33]. Conversely, another study posited that hyperpolarization-activated cation (HCN) channels are responsible for increasing neuronal preference for coherent looming stimuli [20]. The density of HCN channels generally increases with dendritic distance from the soma [53, 54], which suggests that dendritic distance-based synaptic integration in the LGMD dendritic field may also contribute to an increased response to coherent stimuli. Additionally, transient cells have been found to exhibit strong self-inhibition, which may reduce LGMD responses to translating stimuli [55]. Together, these findings indicate that the mechanisms underlying lateral and self-inhibition, as well as excitation in LGMD pathways, are far more complex than our current thoughts. This highlights the need for more investigations into the interplay between these components in generating LGMD’s neuronal responses.
@@
Figure 4: The diagram of a possible visual signal processing of LGMD: there are three key areas represented by question marks that highlight the unknown mechanisms in LGMD circuit processing. First, it remains unclear whether a separate neuron is responsible for gathering global information and subsequently inhibiting other neurons. To date, the exact location for global inhibition is still mysterious. In line with a recent modeling work [22], we assume that this inhibition is mediated by a separate neuron in the lamina layer. Second, the mechanism of lateral excitation remains poorly understood. Research has indicated that lateral excitation operates on a larger time scale compared to lateral inhibition [49], and most likely occurs within trans-medullary afferent neurons (TmAs) [55]. In the processing diagram presented here, we have not explicitly depicted lateral excitation since the neurons in the medulla are classified by ON/OFF contrasts, rather than as specific neurons like TmAs or DUB neurons. Moreover, we assume that ON and OFF signals at the medulla layer are completely separate, as how these signals interact is still unknown. Third, it is uncertain whether interactions occur within the dendritic trees or how integration occurs near the spike initiation zone (SIZ). In this diagram, we assume that the signals from the three dendritic trees are linearly summed without interaction, and then transmitted to the SIZ. After passing through the spike frequency adaptation, the combined signal is subsequently conveyed to the DCMD towards motor system.
@@
Fourthly, dendritic fields B and C of the LGMD1 neuron exhibit phasic ON and OFF inhibition as part of the FFI mechanism, which increases in parallel with the excitation observed in dendritic field A. The FFI signal is dependent on the size of the looming object rather than its approach velocity [12]. Earlier models assumed that FFI was directly received by dendritic fields B and C from TmAs, the same input neurons providing excitation to dendritic field A. However, recent studies have suggested that FFI is derived instead from DUB neurons. The receptive fields of DUB neurons are highly overlapped, and these neurons show no directional selectivity. In general, DUB neuron firing rates persist for at least $100$ ms after the LGMD stops responding and then gradually return to the resting level [56]. Despite these insights, the existence and nature of interactions between DUB neurons (which provide inhibitory input to the LGMD) and TmAs (which provide excitatory input) remain elusive. This gap in knowledge underscores the complexity of inhibitory and excitatory interactions within the LGMD’s afferent circuitry.
@@
From our current understanding, the potential visual signal processing pathway for LGMD can be hypothesized as illustrated in Fig. 4. More specifically, for a looming stimulus, the R cells or photoreceptors within each ommatidium convert the captured luminance changes into electrical signals and pass them to their respective sets of LMCs, where ON/OFF signal separation occurs. Furthermore, global inhibition, which can serve to normalize excitation input to a reasonable level, may occur within each LMC or through a separate pathway. The separated ON/OFF signals are then passed to the medulla layer, where they are processed by TmAs and DUB neurons, respectively. Within the TmAs, self-inhibition reduces the response of the LGMD to translating objects. The TmAs could also exhibit lateral excitation, which enhances LGMD responses to coherent stimuli. The excitatory OFF contrast signals processed by TmAs are then transmitted to dendritic field A of the LGMD, with a portion of the ON contrast excitation signal also passing to dendritic field C. DUB neurons, characterized by relatively large receptive fields, encode the angular size of looming objects in both ON and OFF contrasts and transmit these signals to dendritic fields B and C, respectively. Lateral inhibition occurs at the surface of the LGMD, and synaptic integration can also occur within each dendrite, based on the distance between the dendrite and the soma. Signals from the three dendritic fields are combined at the SIZ, where spike frequency adaptation occurs. This mechanism can be considered a form of habituation, reducing the neuronal response to translating stimuli while enhancing the response to looming stimuli. Finally, the spikes generated by the LGMD are transmitted to its postsynaptic partner DCMD, triggering escape behavior.

### v2

@@
1cm1cm Compared to human vision, locust visual systems excel at rapid and precise collision detection, despite relying on only hundreds of thousands of neurons organized through a few neuropils. This efficiency makes them an attractive model system for developing artificial collision-detecting systems. Specifically, researchers have identified collision-selective neurons in the locust’s optic lobe, called lobula giant movement detectors (LGMDs), which respond specifically to approaching objects.
Research upon LGMD neurons began in the early 1970s. Initially, due to their large size, these neurons were identified as motion detectors, but their role as looming detectors was recognized over time. Since then, progress in neuroscience, computational modeling of LGMD’s visual neural circuits, and LGMD-based robotics have advanced in tandem, each field supporting and driving the others. Today, with a deeper understanding of LGMD neurons, LGMD-based models have significantly improved collision-free navigation in mobile robots including ground and aerial robots.
This review highlights recent developments in LGMD research from the perspectives of neuroscience, computational modeling, and robotics. It emphasizes a biologically plausible research paradigm, where insights from neuroscience inform real-world applications, which would in turn validate and advance neuroscience. With strong support from extensive research and growing application demand, this paradigm has reached a mature stage and demonstrates versatility across different areas of neuroscience research, thereby enhancing our understanding of the interconnections between neuroscience, computational modeling, and robotics. Furthermore, this paradigm would shed light upon the modeling and robotic research into other motion-sensitive neurons or neural circuits.
@@ 1 Introduction
Visually guided collision avoidance is a common behavior in many sighted animal species. Unlike humans, locusts can accurately perceive impending collisions using a neural circuit comprising only hundreds of thousands of neurons, interconnected through a few neuropils. In adult locusts, at least 150,000 neurons contribute to collision detection—a significantly smaller number than invertebrates. This compact neural architecture makes the locust visual system an ideal model for developing efficient and simplified collision detection mechanisms.
@@
Anatomical research on the LGMDs and circuitry was inaugurated in the 1970s [3, 4, 5, 6], while two distinct types of computational models of LGMD1 emerged in the 1990s [7, 8]. The first type of computational models incorporates physical attributes, i.e., image size (angular size) and angular speed on the visual receptive field (RF), to represent the LGMD’s response to looming objects, focusing on dendritic computations within the neuron. In contrast, the second type of models utilizes critical image cues and lateral inhibition within the signal processing pathway to LGMD1 to simulate its neural responses, enabling the distinction between approaching, receding, and translating stimuli. Although LGMD2 shows the similar looming selectivity222neuron responding most strongly to approaching rather than other kinds of visual movements of LGMD1, computational modeling of LGMD2 has emerged late in recent years due to very limited biological findings, compared to LGMD1 [9].
In real-world scenarios, the angular size and speed of an approaching object on the RF are often unknown, unable to measure, which poses significant challenges for implementing the first type of computational model in complex robot scenarios interacting with dynamic environments. Moreover, this model could only predict the LGMD1 response to approaching stimuli, without accounting for other types of motion like receding and translating. In contrast, the second type of model has been successfully extended and demonstrated in real-world environments using a mobile robot system [10], marking a significant milestone in the development of LGMD-based robotics.
@@ 2 How LGMD works as looming detectors
2.1 Biological briefs
@@
Figure 2: The morphology of LGMD1 and LGMD2 with response to different visual stimuli. (A) The afferent network of LGMD1 (adapted from [26]). LGMD1 is located in the fourth layer of the optic lobe in locusts, with the retina, lamina, and medulla serving as its afferent layers. Specifically, there are $\sim 10$ neurons per lamina column with L1 and L2 proposed to be upstream of the LGMD. (B) The morphology of LGMD1 (adapted from [27]). LGMD1 has three distinct dendritic fields where field A receives approximately $\sim 15,000$ excitatory retinotopic inputs from the entire visual hemifield, while fields B and C receive approximately $\sim 500$ non-retinotopic feedforward inhibitory inputs related to ON and OFF contrasts, respectively. Recent research has shown that dendritic field C also contributes to processing excitatory ON-contrast signals. (C) Schematic diagram of LGMD1 and LGMD2 morphology. The LGMD2 neuron (illustrated in blue) has only a single large dendrite field, and its downstream signaling pathway and postsynaptic targets remain unclear. In contrast, the downstream signaling pathway of LGMD1 has been identified as the DCMD. Dendrite field A (depicted in green) receives excitatory inputs and encodes angular velocity in the OFF contrast. The “-” symbol indicates that lateral inhibition may also occur within or preceding the dendrite of LGMD1. Dendrite fields B and C (shown in red) receive inhibitory signals from the ON and OFF pathways, respectively. Additionally, recent studies suggest that dendrite field C also processes non-retinotopic excitatory signals from the ON pathway [16]. (D) The response of LGMD1 when a locust views a square projected on a screen approaching and then receding at $5m/s$ (adapted from [28]). LGMD1 shows a strong and continuous spike train as the square approaches, whereas the neuron displays only a phasic spike at the onset of receding. (E) The response of LGMD1 to approaching and translating stimuli (adapted from [19]) - regardless of the size and speed of the translating bar, the instantaneous firing rate of LGMD1 is significantly lower when the locust views translating stimuli compared to looming stimuli. (F) The neuronal responses of LGMD1 and LGMD2 when the locust views light and dark rectangles approaching and receding, as well as in response to changes in luminance (adapted from [9]) - LGMD2 is selectively excited by darker objects approaching. The long arrow indicates LGMD2 excitation after the onset of light object receding, while the short arrow indicates hyperpolarization of LGMD2 when both light and dark objects ceased during approach or at the start of receding of a dark object.
@@
2.2 Classic computational theory on LGMD
Regarding computational modeling of LGMDs, the models of LGMD1 emerged since the 1990s [8], while the first model of LGMD2 was proposed over two decades later [31, 32]. In general, bio-inspired looming perception visual systems can be categorized into two primary classes. The first class focuses on mapping nonlinear relationships between the physical attributes of stimuli (image size and image velocity) and corresponding neuronal activity. The second class utilizes hierarchical neural networks modeled after well-established anatomical pathways responsible for looming perception.
Assuming that dendritic field A of the LGMD1 neuron encodes angular velocity, while dendritic fields B and C encode angular size, Hatsopoulos et al. proposed the first computational model of LGMD1, known as the $\eta$-function [7]. As illustrated in Fig. 3A, for a looming object with half-size $l$ and constant approach speed $v$, the angular size $\theta$ projected on the retina can be described by $\tan(\theta)=\frac{2l}{v\cdot t}$, which increases rapidly as collision becomes imminent. The angular velocity $\theta^{\prime}(t)$ is the derivative of angular size with respect to time. The neuronal response of LGMD1 to looming stimuli, from a single neuron computational perspective, is modeled by the following $\eta$-function [^3] Responses of looming sensitive neurons in both vertebrates and invertebrates could also encode other optical variables like $\tau$ relative rate of expansion, $\rho$ absolute rate of expansion [33].:
@@ 2.2 Classic computational theory on LGMD
Researchers have demonstrated that the multiplicative combination of neuronal signals encoding an object’s angular size and angular speed can be achieved through a logarithmic-exponential transformation in the LGMD1 neuron [38], providing further evidence that the mathematically straightforward $\eta$-function can serves as the intrinsic computational mechanism for the neuronal response of LGMD1. Initially, the $\eta$-function models behave like the LGMD1 only for objects approaching at a constant speed [39], lacking the capacity to simulate detailed responses to accelerating objects and other types of visual stimuli, such as recession and translation. Recently, Dewell et al. demonstrated that LGMD responses to approaching objects with non-constant velocities are accurately predicted by the $\eta$ model, exhibiting only minor timing discrepancies [40].
On the other hand, Rind and Simmons proposed a four-layered neural network to model the LGMD1’s neuronal response to looming, receding, and translating stimuli [8] (see Fig. 3C). In this network, critical image cues (denoted as P in Fig. 3C) are generated by edges that change in extent or velocity during object approach, resulting in strong excitation of computational units (denoted as E in Fig. 3C). Lateral inhibition within these units is propagated with a time delay (denoted as I in Fig. 3C). An additional unit, termed the F unit (denoted as F in Fig. 3C), is integrated into the network to simulate the feed-forward inhibition (FFI) observed in the visual pathway to LGMD1. The captured critical image cues, the critical race between excitation and inhibition and the FFI together fine tune the looming selectivity of the proposed neural network, showing strong and increasing response to looming stimuli, while transient response to object recession and minor response for single edge crossing (elongation) (see Fig. 3D, E). Reducing the strength of lateral inhibition significantly raises the network’s response to object receding and translating. Moreover, removing the FFI prolongs the network’s response to both object approach and recession.
@@ 2.3 Controversial arguments on LGMD neural processing
Although both categories of computational models effectively capture the neuronal characteristics of the LGMD1 when faced with looming objects, a key point of contention between these models lies in the timing of the peak response. In the first class of models, the $\eta$-function produces a peak response after a fixed delay, following the point at which the looming object reaches a specific image size threshold. However, the peak will be occurring after the object projects its maximum size onto the retina when $l/|v|\leq 5\text{ ms}$ [42]. This ratio can be achieved by, for example, an 8 cm-wide object moving at $8$ m/s ($28.8$ km/h) or a 4 cm-wide object traveling at $4$ m/s ($14.4$ km/h). In contrast, some experimental results for the second category of models, as reviewed in [23], indicate that the peak response consistently occurs before the looming object reaches its maximum size on the RF.
@@
Secondly, while ON and OFF contrasts are separated in lamina monopolar cells (LMCs) located in the lamina, the precise mechanisms of how these signals combine or interact, particularly in the optic lobe of locusts, are unverified. Earlier research on LGMD1 suggested that ON and OFF contrast signals jointly flow into the medulla layer, where they transmit to dendritic field A for excitation, and to dendritic trees B and C for ON/OFF-contrast FFI, respectively [37]. However, recent neuroscience studies indicate that dendritic field C may partially contribute to excitation in response to ON contrast stimuli by receiving inputs from neurons in the dorsal uncrossed bundle (DUB), which possess smaller receptive fields ($\sim 10\degree$). In contrast, OFF contrast-sensitive neurons responsible for FFI are fewer in number and exhibit broader receptive fields ($\sim 50\degree$) [20, 16, 50]. Moreover, it has been mentioned that global inhibition, scaled by overall laminar activity, functions as a normalization mechanism in the medulla layer for increasing excitation as collision approaches [51]. Computational modelers have attempted to replicate this function by introducing a non-spiking neuron in the lamina layer [22]. Such GI neuron gather signals across LMCs then affect the trans-medullary afferent neurons (TmAs) in the medulla where the authors suggested it may function as a network of neurons. In addition, they also modeled lateral inhibition crossing between units directly from the LMCs [22]. However, anatomical evidence may conflict with this modeling assumption, as each ommatidium has its own set of photoreceptors and LMCs, and LMCs are believed to interact only within their own set and not across different ommatidia [52].
@@
Thirdly, regarding the TmAs as inputs to LGMDs (targeting dendritic field A of both LGMD1 and LGMD2), early neural modeling efforts assumed that TmAs were responsible for lateral inhibition in the LGMD pathway [21, 22]. When traced for reconstruction, TmAs indeed exhibited lateral connections and received input from the outer medulla layer. Most works demonstrated such lateral interactions are likely inhibitory types [53, 54, 17]. However, evidence suggests that this lateral interaction contributes to excitation rather than inhibition[51], and the lateral excitation involving TmAs is thought to enhance LGMD’s responses to coherent stimuli [51]. Another study posited that hyperpolarization-activated cation (HCN) channels are responsible for increasing neuronal preference for coherent looming stimuli [20]. Through two works specializing in dendritic synaptic integration [55, 56], we infer that the density of HCN channels may generally increase with dendritic distance from the soma, i.e., distance-based synaptic integration in the LGMD’s dendritic field may also contribute to an increased response to coherent stimuli. Additionally, transient cells may exhibit strong self-inhibition, which could reduce LGMD responses to translating stimuli [57]. Together, these findings indicate that the mechanisms underlying lateral and self-inhibition, as well as excitation in LGMD pathways, are far more complex than our current thoughts. This highlights the need for more investigations into the interplay between these components in generating LGMD’s neuronal responses.
@@
Figure 4: The diagram of a possible visual signal processing of LGMD from our perspective and assumption: there are three key areas represented by question marks that highlight the unknown mechanisms in LGMD circuit processing. First, it remains unclear whether a separate neuron is responsible for gathering global information and subsequently inhibiting other neurons. To date, the exact location for global inhibition is still mysterious. In line with a recent modeling work [22], we assume that this inhibition is mediated by a separate neuron in the lamina layer. Second, the mechanism of lateral excitation remains poorly understood. Research has indicated that lateral excitation operates on a larger time scale compared to lateral inhibition [51], and most likely occurs within trans-medullary afferent neurons (TmAs) [57]. In the processing diagram presented here, we have not explicitly depicted lateral excitation since the neurons in the medulla are classified by ON/OFF contrasts, rather than as specific neurons like TmAs or DUB neurons. Moreover, we assume that ON and OFF signals at the medulla layer are completely separate, as how these signals interact is still unknown. Third, it is uncertain whether interactions occur within the dendritic trees or how integration occurs near the spike initiation zone (SIZ). In this diagram, we assume that the signals from the three dendritic trees are linearly summed without interaction, and then transmitted to the SIZ. After passing through the spike frequency adaptation, the combined signal is subsequently conveyed to the DCMD towards motor system.
@@
Fourthly, dendritic fields B and C of the LGMD1 neuron exhibit phasic ON and OFF inhibition as part of the FFI mechanism, which increases in parallel with the excitation observed in dendritic field A. The FFI signal is dependent on the size of the looming object rather than its approach velocity [12, 50]. Earlier models assumed that FFI was directly received by dendritic fields B and C from TmAs, the same input neurons providing excitation to dendritic field A. However, recent studies have suggested that FFI, delivered to field C, is derived instead from DUB neurons. The receptive fields of DUB neurons are highly overlapped, and these neurons show no directional selectivity. In general, DUB neuron firing rates persist for at least $100$ ms after the LGMD stops responding and then gradually return to the resting level [50]. Despite these insights, the existence and nature of interactions between DUB neurons (which provide inhibitory input to the LGMD) and TmAs (which provide excitatory input) remain elusive. This gap in knowledge underscores the complexity of inhibitory and excitatory interactions within the LGMD’s afferent circuitry.
@@
From our current understanding, the potential visual signal processing pathway for LGMD can be hypothesized as illustrated in Fig. 4. More specifically, for a looming stimulus, the R cells or photoreceptors within each ommatidium convert the captured luminance changes into electrical signals and pass them to their respective sets of LMCs, where ON/OFF signal separation occurs. Furthermore, global inhibition, which can serve to normalize excitation input to a reasonable level, may occur within each LMC or through a separate pathway. The separated ON/OFF signals are then passed to the medulla layer, where they are processed by TmAs and DUB neurons, respectively. Within the TmAs, self-inhibition reduces the response of the LGMD to translating objects. The TmAs could also exhibit lateral excitation, which enhances LGMD responses to coherent stimuli. The excitatory OFF contrast signals processed by TmAs are then transmitted to dendritic field A of the LGMD, with a portion of the ON contrast excitation signal also passing to dendritic field C. DUB neurons, characterized by relatively large receptive fields, encode the angular size of looming objects in both ON and OFF contrasts and transmit these signals to dendritic fields B and C, respectively. Lateral inhibition occurs at the surface of the LGMD, and synaptic integration can also occur within each dendrite, based on the distance between the dendrite and the soma. Signals from the three dendritic fields are combined at the SIZ, where spike frequency adaptation occurs. This mechanism can be considered a form of habituation, reducing the neuronal response to translating stimuli while enhancing the response to looming stimuli. Finally, the spikes generated by the LGMD are transmitted to its postsynaptic partner DCMD, triggering escape behavior. In general, this hypothesized model could replicate biological functionality, though certain implementation details do not precisely reflect biological realism.